OpenStudy (anonymous):

PLEASE, someone help! 1. What is the equation of the line which includes points (1, 3) and (2, 8)?

OpenStudy (anonymous):

\[m=y _{2}-y _{1}/x _{2}-x_{1}=8-3/2-1=5\]

OpenStudy (anonymous):

You basically use Y2 - Y1/ X2 - X1 to get the slope. That case it'll be 5/1 so the slope will be 5. Then it's y = mx + b. You already know that the slope is 5 so it will be y = 5x + b. Then you plug in one of those points to find the y intercept. 8-3 / 2 - 1 = 5 Y = 5x + b, plug in (2,8) 8 = 5(2) + b 8 - 10 = b b = -2 therefore y = 5x - 2

OpenStudy (anonymous):

then use any point in \[y-y_{1}=m(x-x _{1})\]

OpenStudy (anonymous):

ie \[y-3=5(x-1)\] then solve

OpenStudy (anonymous):

thats correct do you understand the logic